Diffusion Morning Report 04/07/06

Ted Berg tberg at snf.stanford.edu
Fri Apr 7 06:39:09 PDT 2006


Good Morning All,
    Looking good this morning. Epi is  doing Ok . TEOS seems to be 
depositing correctly. Tylan 3 has bad uniformity . Tylan 1 has temp
oscillations. Everything else looks OK for now. Burn box cleaned this 
morning.


	Date: 	4/7/06 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	
	
	
Tylan 1 	Up 	  	
	
	
Tylan 2 	Up 	  	
	
	
Tylan 3 	Up 	  	
	
	
Tylan 4 	Up 	  	
	
	
Tylan 5-BBr3 	Up 	  	
	
	
Tylan 6-POCl3 	Down 	temp cal 	
	
	
Tylan 7-TEOS2 	Up 	  	16.69/30 	
	
Tylan 8-SiGe 	Up 	  	26.82/30 	
	
Tylan 9-Poly 	Up 	  	17.66/30 	
	
Tylan 10-Nitride 	Up 	  	18.35/30 	
	
Tylan 11-BPSG-LTO 	Up 	  	16.35/30 	
	
Tylan 12-FGA 	Up 	  	  	
	
  	  	
	
	
	
Thermco 1 	Up 	  	
	
	
Thermco 2 	Up 	  	
	
	
Thermco Poly 	Up 	  	
	
	
Thermco Nitride 	Down 	Testing 	
	
	

	
	
	
	
	
Epi 	up 	  	
	
	

	
	
	
	
	
Burn Box 	Up 	 Cleaned this morning
	
	
	
ZYGO 	up 	  	   	
	

	
	
	
	
	
GAAS 22 	UP 	
	
	
	
GAAS 23 	up 	  	
	
	

	
	
	
	
	
SiH4 for furnaces 	
	A on line 645 psi 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060407/27c51451/attachment.html>


More information about the diffusion mailing list